OpenStudy (anonymous):

What are the lengths of the legs of a right triangle in which one acute angle measures 19° and the hypotenuse is 15 units long?

OpenStudy (anonymous):



OpenStudy (anonymous):

sin 19=p/h .14=p/15 p=2.1

OpenStudy (anonymous):

base=14.8